A basketball league has 90 games in a season. How many teams are there if each team plays every team twice?
Answer by venugopalramana(3286)

A basketball league has 90 games in a season. How many teams are there if each team plays every team twice?
LET THE NUMBER OF TEAMS =N
NUMBER OF MATHCHES IN 1 ROUND = NC2=N(N-1)/2
NUMBER OF MATCHES IN 2 ROUNDS = 2N(N-1)/2 = 90
N^2-N-90=0
N^2-10N+9N-90=0
N(N-10)+9(N-10)=0
(N-10)(N+9)=0
N=10